A plane flies 440 miles with wind and 340 miles against wind in the same length of time. If the speed of the wind is 25 mph, find the speed of the plane in still air.
Answer by Alan3354(69443)

Similar problem.
================================
A boat can travel 39 mph in still water.If it travels 235 miles with the current in the same length of time it travels 155 miles against the current, what is the speed of the current?
----------
Time upstream = 155/(39 - c)
Time downstream = 235/(39+c)
============
155/(39 - c) = 235/(39+c)
31/(39 - c) = 47/(39+c)
31*(39 + c) = 47*(39 - c)
etc